If the hydroxide concentration of a CH3COOH/CH3COO- buffer system is 0.000045 M, what is the acetate...

  1. If the hydroxide concentration of a CH3COOH/CH3COO- buffer system is 0.000045 M, what is the acetate concentration if the acetic acid concentration is 0.74 and the dissociation constant, Ka, of CH3COOH is 6 x 10-9 ?

  1. A hot piece of metal at 450K is placed in pure water that has a volume of 55.0 mL which has an initial temperature of 25.3°C. After five minutes, equilibrium is established and a final temperature of 44.4°C is reached . What is the mass of the metal, if the metal has a specific heat 0.33 J/g°C
  1. An unlimited amount of potassium permanganate reacts with a 10.0 mL solution of 0.1 M of 7-fluoro-5-isopropyl-2-octene. How many grams of organic compound product can be produced?
Answer #1

Ans 1

[OH-] = 0.000045 M

pOH = - log [OH-] = - log (0.000045) = 4.3468

pH = 14 - pOH

= 14 - 4.3468

= 9.6532

Ka = 6*10^-9

pKa = - logKa = - log(6*10^-9) = 8.222

From the Henderson Hassalbalch equation

pH = pKa + log([acetate] / [acetic acid])

9.6532 = 8.222 + log ([acetate] / [0.74])

1.4312 = log ([acetate] / [0.74])

26.989 = [ acetate] / [0.74]

[ acetate] = 19.972 M